Translate the problem into a pair of linear equations in two vaiables. Solve the equations using either elimination or substitution. State your answer for the specified variable.
Andy has 14 coins made up of quarters and half-dollars, and their total value is $4.25. How many quarters does he have?

Found 2 solutions by richwmiller, solver91311:
Answer by richwmiller(17219)

let q=quarters and h=half dollars
q+h=14
q=14-h
25q+50h=425
25(q+2h)=25*17
divide by 25
q+2h=17
14-h+2h=17
14+h=17
3=h
q=11
check
11*25+3*50=425
10*25=250
11*25=275
275+150=425
425=425
ok


Answer by solver91311(24713)


Let represent the number of quarters. Let represent the number of half-dollars. Then is the value of the quarters in cents and is the value of the half-dollars in cents. The total amount of money he has, expressed in cents, is 425.

He has 14 coins:



The total value is $4.25:



Either of the elimination or substitution will work just fine to solve this system, but given the simplicity of the first equation, I would choose the substitution method.

Let me know what you obtained for an answer.

John